OpenStudy (anonymous):

Evaluate x 1 - x -1 + x 0 for x = 2.???

OpenStudy (anonymous):

is the question "evaluate x: 1-x-1+x=0, for x =2?"

OpenStudy (anonymous):

its: Evaluate x ^1 - x ^-1 + x^0 for x = 2.

OpenStudy (anonymous):

\[x-{ 1/x }+1=2-1/2+1=5/2\]